Plot the point whose polar coordinates are given. Then find the Cartesian coordinates of the point. (a) (2, 34/2)
(x, y) = ( D (b) (2V2, A/4)
(x, y) - ( (c) (-9, -A/6) --8 -6 -4 - 46
Answer #1

polas Coordinate = cartesian Coordinate (0,0) (ny) x=8coso T = & sind @ _ [2, 38/31 maa Q = 3 1/2 cose, a Cos ( 381 ) = 0 22(-9, X16) Q₂ 2216 X = r costs = -9X1322 2 3 7 2 8 gina/6 a 9x1 - 96 So; cartesian point = (-939 LE(-7.794 45 8 65 ---2-1 tu +